# Frieren: Efficient Video-to-Audio Generation with Rectified Flow Matching

### v1

@@ 3.3 Re-weighting RFM objective with logit-normal coefficient
We re-weight the RFM objective with this weighting function to replace the original objective.
@@ 4.3.1 Model size of vector field estimator
We observe that when the model parameters are reduced from about 159M to 71M, performance declines across all metrics, where FD, KL, FAD, and KID increase by 0.77, 0.05, 0.18, and $\times 10^{-3}$, and IS, alignment accuracy, MOS-Q and MOS-A drop by 0.26, 1.18%, 0.07 and 0.07, respectively. On the other hand, when the parameter number increases to 421M, the model obtains a performance improvement across multiple metrics, with FD, KL, FAD, and KID declining by 0.06, 0.05, 0.02, and 0.08×$\times 10^{-3}$, and IS, alignment accuracy, and MOS-Q increasing by 0.03, 0.09%, and 0.03. This result aligns with our common sense that a larger model capacity leads to better performance. Moreover, it indicates the potential scalability of our model to large model size and amount of data, which may bring it broader prospects for applications. Details of model parameters are provided in appendix A.
@@ 4.3.2 Visual feature characteristics
In table 4, we compare the results of Frieren using two different types of visual features from CAVP and MAViL. Intuitively, the MAViL feature should be more robust and contain richer audio-related semantic information, as it utilizes masked-reconstruction together with inter-modal and intra-modal contrastive learning, in contrast to CAVP trained solely with inter-modal contrastive learning. On the other hand, however, due to MAViL’s convolutional downsampling in the temporal dimension, its feature sequence has a lower effective FPS of 2 with the same 4 FPS video input as CAVP. The results in the table indicate that the model with MAViL feature excels in audio quality and diversity, with differences of FD, IS, KL, and FAD being 0.17, 0.25, 0.24, and 0.06. Meanwhile, it exhibits a 7.05% decrease in alignment accuracy. This result yields two insights for V2A tasks: 1) at relatively low frame rates, the frame rate of features, rather than content, is more likely to become the bottleneck for generating visual-synchronized audio; 2) compared to high video frame rates, the semantic information and robustness of visual features are more crucial for the quality of generated audio.
@@ 4.3.4 Re-weighting RFM objective
We conduct ablation on RFM objective re-weighting and report the results in table 5. We can see that discarding objective re-weighting leads to a comprehensive deterioration in FD, IS, KL, FAD, and alignment accuracy, with differences being 0.13, 0.22, 0.01, and 0.07. It also leads to a drop of MOS-Q and MOS-A. This validates the positive impact on model performance of objective re-weighting as an equivalent operation of adjusting sampling probabilities at different time positions.

### v2

@@ 3.3 Re-weighting RFM objective with logit-normal coefficient
We re-weight the RFM objective with this weighting function to replace the original objective and observe in our experiment that this re-weighting helps to slightly improve audio quality and temporal alignment at the cost of a marginal decrease in audio diversity.
@@ 4.3.1 Model size of vector field estimator
We observe that when the model parameters are reduced to 71M, performance declines across all metrics, where FD, KL, FAD, and KID increase by 0.77, 0.05, 0.18, and $\times 10^{-3}$, and IS, alignment accuracy, MOS-Q and MOS-A drop by 0.26, 1.18%, 0.07 and 0.07, respectively. However, when the parameter number increases to 421M, there is a performance degradation across multiple metrics, with KL, FAD, and KID increasing by 0.03, 0.04, and 0.48×$\times 10^{-3}$, and IS, alignment acc declining by 0.13 and 2.06%. We speculate that this anomalous phenomenon may be due to the convergence difficulty for the larger model under similar training steps, or the redundant model capacity tends to cause overfitting on a relatively small dataset like VGGSound, deteriorating the model’s generalization performance. In summary, we achieve relatively balanced model performance with the parameter of the estimator being around 160M. Details of model parameters are provided in appendix A.
@@ 4.3.2 Visual feature characteristics
In table 4, we compare the results of Frieren using two different types of visual features from CAVP and MAViL. Intuitively, the MAViL feature should be more robust and contain richer audio-related semantic information, as it utilizes masked-reconstruction together with inter-modal and intra-modal contrastive learning, in contrast to CAVP trained solely with inter-modal contrastive learning. On the other hand, however, due to MAViL’s convolutional downsampling in the temporal dimension, its feature sequence has a lower effective FPS of 2 with the same 4 FPS video input as CAVP. The results in the table indicate that the model with MAViL feature excels in audio diversity, with differences of FD, KL, and FAD being 0.17, 0.24, and 0.06. Meanwhile, it exhibits a 7.05% decrease in alignment accuracy and a 0.25 decrease in IS. This result yields two insights for V2A tasks: 1) at relatively low frame rates, the frame rate of features, rather than content, is more likely to become the bottleneck for audio quality and visual-audio synchrony; 2) compared to high video frame rates, the semantic information and robustness of visual features are more crucial for the diversity of generated audio.
@@ 4.3.4 Re-weighting RFM objective
We conduct ablation on RFM objective re-weighting and report the results in table 5. We can see that compared to the vanilla objective, introducing re-weighting results in improvements of 0.22 and 0.18% for IS and alignment accuracy. This validates the positive impact of objective re-weighting on audio quality and temporal alignment. On the other hand, objective re-weighting causes a decrease in audio diversity, with differences in FD, FAD, and KID being 0.30, 0.07, and 0.37×$\times 10^{-3}$, respectively.